OpenStudy (anonymous):

The area of a rectangle is 61.64 square feet and the length is 9.2 feet. What is the width of this rectangle?

OpenStudy (anonymous):

6.7 feet

OpenStudy (anonymous):

@lol23 @Fullerene

OpenStudy (anonymous):

A=l*w 61.64=9.2*w w=6.7ft

OpenStudy (anonymous):

\[61.64 \div 9.2\]